Please help me solve this equation 5n^2-10n+7=0
Answer by stanbon(75887)

solve this equation 5n^2-10n+7=0
-----------------
Use the Quadratic Formula::
n = [10 +- sqrt(100-4*5*7)]/10
----
n = [10 +- sqrt(-40)]/10
----
n = [10 +- 2isqrt(10)]/10
-----
n = 1+(1/5)sqrt(10)i or n = 1-(1/5)sqrt(10)i
=================
Cheers,
Stan H.
=================